OpenStudy (anonymous):

Charla and David are observing the leaf growth rates of their plants. Charla’s plant starts on day 1 with 43 leaves and gets 1 new leaf a day. David’s plant starts on day 1 with 19 leaves and gets 4 leaves a day. How many days will it take David’s plant to have as many leaves as Charla’s plant? 7 5 9 8

thomaster (thomaster):

what is the variable and what is the constant for each plant?

OpenStudy (anonymous):

For this question, you could do it by just trial and error. If you want to mathematical approach... Pc(Charla's plant)=42+n (where n is the number of days) Pd = 7n so we want to find when Pd=Pc Solving.. 7n=42+n 6n=42 n=7. So, on day 7, David's plant will have as many leaves as Charla's plant.

OpenStudy (anonymous):

Therefor answer is the first one. i hope i helped you. if you think i did please mark me as you "best answer" it will be greatly appreciated if you do.

thomaster (thomaster):

@Gio123 where did you get 7n? The part for charla's plant is correct. n+43 The part for David's plant should be 4n+19 n+43=4n+19 43=3n+19 3n=24 n=8